### 18 Dec 2023 v1

@@ 2.2 Overview of vision-based seizure semiology analysis
Seizures characterized by motor manifestations are analyzed and classified based on the specific type of visible (mainly motor) symptomatology. Simple motor seizures are defined by unnatural movements, and are categorized into myoclonic, clonic, tonic, versive, and tonic-clonic seizures. This classification depends on factors such as the duration of muscle contraction, rhythmicity of movement repetition, and the muscle groups or body segments involved (e.g. asymmetric posture, flexion of the neck, abduction of both arms, and turning of eyes and head to one side). In contrast, more complex motor manifestations may involve complex and sometimes seem naturalistic, and comprise action patterns involving movements across different body segments (e.g. manual and oral automatisms such as chewing, swallowing, smacking the lips and fumbling, e.g. with clothes) [149, 30]. Some existing research on automated quantification of semiology is based on the hypothesis that seizures with similar motor symptoms may involve neuronal activity within the same specific brain networks, and thus could be sufficient to detect and categorize patients with specific types and brain localizations of epilepsy [113]. Vision-based approaches for semiology highlight the significance of analyzing spatio-temporal information during seizures, employing action recognition methodologies. In this domain, the discriminative power of deep learning architectures is exploited for markerless motion capture (MoCap) of body regions for isolated clinical manifestations from the body, head, face, and hands. Furthermore, deep learning is integral to seizure detection (identifying epileptic events within a continuous data stream) and seizure classification (categorizing seizures or seizure disorders).
@@
Deep learning models play a crucial role in advancing motion capture technologies, particularly for detecting regions of interest (ROI). These ROIs may encapsulate entities like patients, heads, faces, or hands. Yet these techniques go beyond mere object identification and localization, inferring semantic anatomical landmarks such as facial landmarks, body keypoints, and hand landmarks from video data. While seizures, strictly speaking, may not be categorized as actions, HAR models are designed to encode robust representations of human motion that may prove relevant for seizure characterization. These methods are commonly structured to classify human actions by aggregating predictions for snippets sampled from short clips. Traditionally, a subject-level prediction is obtained by averaging all predictions made at the snippet level.
The deep learning methods may also be trained in different manners. In a supervised learning paradigm, training occurs offline as the machine utilizes labeled data to predict outcomes of unforeseen data. However, the scarcity of labeled clinical data and challenges in acquiring additional data make efficient data utilization crucial. To overcome the limitations posed by small datasets in Epilepsy, motion capture and HAR architectures rely on transfer learning. Networks are pre-trained on large-scale datasets [76] which allows the architecture to learn a robust general representation, which can then be applied to extract features and learn a representation specifically for seizure video recordings. However, the generalizability of trained classifiers in subject-independent classification settings is hindered by the considerable variation in behavior across different subjects. A research direction with significant potential is self-supervised learning. In scenarios where no class labels are available, an effective representation could potentially be learned through an end-to-end method entirely in an unsupervised manner, utilizing a large volume of unlabeled semiology data.
@@
Figure 1: Overview of deep learning applications for semiology analysis. A. Motion capture and skeleton-based human action recognition. B. Spatio-temporal human action recognition. C. Feature representations of the motion exhibited by a patient during a seizure, which is used to identify if new data samples are similar to known semiologies, enabling the categorization of epilepsy types.
@@
In the pre-deep learning era, various hand-crafted feature-based approaches, such as space-time volume-based methods, space-time interest point-based methods, and trajectory-based methods, were employed for RGB video-based detection and HAR. However, with the advent of deep learning, contemporary research in semiology has shifted to the adoption of different types of deep learning frameworks. These frameworks fall into two categories, skeleton-based and spatio-temporal-based HAR. A visual representation of these frameworks is provided in Fig. 1. The first, skeleton-based HAR, focuses on detecting semantic key points or movement direction/intensity initially. Subsequently, quantitative kinematic and spectral parameters are computed to detect/classify epilepsy types with classical machine learning or deep learning methods. Here, skeleton information refers to body, facial, or hand landmarks. These approaches have been utilized by [10, 9, 94] (See Fig. 1A). The second framework, a spatio-temporal-based HAR (frame-based), captures the dynamic variation of the body’s physical appearance from sequences of images. This approach considers both the presence of specific movements of interest (MOI) and their dynamics and biomechanical characteristics (e.g. speed/acceleration patterns, movement amplitude). Works including  [8, 104, 162, 227, 88] are example of this approach (See Fig. 1B). Seizure semiology encompasses the study of the stepwise/temporal progression of a constellation of signs that are reflective of dynamic changes within connected neuronal networks. Thus, investigating a single clinical sign in isolation is generally less informative than considering all signs jointly. As such, multi-stream frameworks have been proposed, capable of jointly quantifying semiologic features from different body locations through fusion or hierarchical approaches (e.g. an ensemble model learning from pose and face skeleton-based streams). Hybrid networks, which fuse the benefits of both skeleton-based and spatio-temporal approaches for seizure detection and classification, have also been proposed in [4, 87, 89]. Characteristic combinations of clinical features of different seizure types can be demonstrated to be associated with seizures arising from different brain regions at group level. For some types of epilepsy, it is the motor features of the seizures that seem to most strongly predict brain localization [32]. Such seizure types lend themselves well to analysis using computer vision-based approaches. A visual example of these motions is illustrated in Fig. 1C. The utilization of these feature representations contributes to the development of robust models capable of discerning and categorizing diverse seizure types. Further details on recent case studies of vision-based approaches in Epilepsy are discussed in Section 3. The following subsections provide details of the specific techniques used by each human action recognition framework.
@@ 2.3 Skeleton-based human action recognition
In skeleton-based action recognition, the primary focus is on explicitly tracking and analyzing the movement and positions of skeletal joints from the body, face, or hands, that inherently contain spatiotemporal information. Skeleton data offers numerous advantages for HAR including a simple yet informative representation, scale invariance, and robustness to variations in clothing textures and backgrounds. These approaches are often fast since the human pose representation is very compact. However, they lack appearance and detailed shape information and the underlying skeleton representation can be considered a noisy representation [189]. Another drawback is the propagation of errors from the motion capture to the action recognition stage, a factor highly dependent on the quality of the spatio-temporal feature extraction of the initial stage. Joint information is vulnerable to conditions where joint estimation is performed erroneously, which is not uncommon due to the frequent occlusions that occur in seizure videos. Fig. 2 illustrates these stages.
@@ 2.3.1 Markerless motion capture: keypoint detection
Human pose estimation: Human pose estimation involves localizing anatomical keypoints, and is widely studied in computer vision and has broad applications (e.g. motion analysis, healthcare, virtual reality). Although camera-based methods for human pose estimation have been studied extensively, in-bed pose analysis comes with specific challenges including joints being occluded, blur, and low-light conditions. Traditional pose estimation models are broadly categorized into two groups: in the first approach, a human bounding box is initially detected, and then keypoints indicating joint locations for each person are estimated (known as top-down methods, e.g. AlphaPose model[63], ViTPose [221]). This can lead to errors as the predicted bounding box in the first step may not include the required joints for the second step. In the second approach, all keypoints are detected in the first stage and they are assigned to each person in an image in a second stage (bottom-up methods, or box-free human detection, e.g. OpenPose  [34]). Human pose estimation has also been investigated in 3D space to predict the locations of body joints. For instance,  [193] developed a 3D human pose estimation algorithm by taking into account spatial and temporal information. In addition to estimating the 3D pose, some methods also recover a 3D human mesh from images or videos [130, 156]. This field has attracted much interest in recent years since it is used to provide extensive 3D structure information related to the human body.
@@
Hand landmark estimation: Hand pose estimation refers to methods used to detect hand landmarks or keypoints. Methods that can localize hand joints share several properties with human body pose estimation, and many approaches proposed for the human body can be adapted for hand pose estimation. Similar to body pose, hand pose can also be estimated in 2D [180, 210] or 3D [153, 46] space. Image generation and synthetic data have also become popular in hand pose estimation. Examples of recently introduced synthetic data are  [109], which includes multi-view hand data, and [122], which features 1 million photo-realistic images of hands interacting with various objects, backgrounds, and textures. For more information about 2D/3D or hybrid hand pose models, refer to [45, 150].
@@ 2.3.2 Action recognition models (skeleton-based HAR)
CNN-based methods: This approach utilizes the advantages of CNN-based object recognition in a 2D space (image space). It maps a 3D representation of the 3D human skeletons into a 2D array (possibly capturing the spatial relationships between the skeleton joints) to learn spatial-temporal skeleton features. Examples of recent works using CNN-based spatiotemporal approaches are CNN fusion  [124], and PoseConv3D [59]. A CNN fusion model for skeletal action recognition is proposed in [124], which uses spatiotemporal information including skeletal trajectory shape images as well as skeletal pose image sequences. PoseConv3D [59] used the 3D-CNN to capture the spatio-temporal dynamics of skeleton sequences; where the input to the 3D-CNN backbone are 3D heatmap volumes. The pseudo heatmaps for joints and limbs that are generated provide rich information as inputs for 3D-CNNs. More information about convolutional neural network-based action can be found in [228].
@@
GNN-based methods: Early approaches to skeleton-based HAR employed RNNs and CNNs, but recent advancements have surpassed them with the introduction of Graph Convolutional Networks (GCNs). Leveraging the expressive power of graph structures, GCNs efficiently represent non-Euclidean data, allowing them to capture both spatial (intra-frame) and temporal (inter-frame) information. These learning models have been adopted for various tasks including skeleton-based human action recognition [178, 152] and facial landmark-based emotion recognition [148]. The spatial-temporal graph (ST-GCN) [224], the first work utilizing GCNs for HAR, is constructed such that the graph nodes represent the joint and the edges encode natural connections in both human body and time. Multiple layers of spatial-temporal graph convolution operations are applied to the input data, generating higher-level feature maps on the graph, subsequently classifying the input into the corresponding action category. Followed by the success of ST-GCN, actional-structural graphs have been proposed to capture both actional links and structural links over time [123]. Another work involves using a graph network to perform spatial reasoning and temporal stack learning (SR-TSL) for skeleton-based HAR, comprising a spatial reasoning network (SRN) and a temporal stack learning network (TSLN) [179]. More recent approaches include ST-GCN++ [58] and Efficient GCN [186].
@@
Transformer-based methods: Vision transformers are emerging as a powerful tool to solve computer vision problems. Recent techniques have also proven the efficacy of transformers beyond the image domain, extending to diverse video-related tasks, including action recognition [17]. The architecture of transformers, featuring a self-attention mechanism, proves useful for estimating temporal relationships and motion dynamics encoded in joint positions. Some examples of these approaches include PoseConv3D [59], SkeletonMAE [218] and InfoGCN++ [47]. PoseConv3D [59] addresses limitations in current Skeleton-based Human Activity Recognition (HAR) methods by combining 3D CNN and a transformer-based approach for feature extraction, modeling temporal dependencies, and handling the sparsity of skeleton data. SkeletonMAE [218] focuses on capturing structural information and temporal dependencies in skeleton data by representing it as a graph. It introduces masked autoencoders to learn meaningful representations by reconstructing the original skeleton sequences with masked nodes. Furthermore, InfoGCN++ [47], leverages graph convolutional networks (GCNs) as part of the transformer network to learn informative representations by predicting future frames in the skeleton sequence. These transformer-based methods incorporate temporal information through a joint learning strategy, resulting in improved performance in action recognition tasks.
2.4 Spatio-temporal-based HAR modality
@@
In "spatio-temporal-based" (or "frame-based") action recognition, the focus is on analyzing regions or patches within sequences of images. This approach does not rely on explicit skeletal data but rather on the visual information captured in the frames, including the spatial layout and temporal changes in these regions [189]. This framework is illustrated in Fig. 3. As these approaches directly operate on RGB videos, there is a possibility of privacy leakage of sensitive patient data from videos. Moreover, obtaining consent from patients to share their raw RGB video data for inter-cohort validation studies and developing approaches that generalize to large populations becomes challenging.
@@
Figure 3: Overview of spatio-temporal human action recognition. A. Motion capture: Region of interest detection such as the body, face, or hands. B. Spatio-temporal representations are learned from sequential frames. CNN-RNN-based, 3D-CNN-based and Transformer-based methods are all commonly used with this type of approach.
@@ 2.4.1 Markerless motion capture: region of interest detection
2.4.2 Action recognition models (spatio-temporal-based HAR)
Spatio-temporal human action recognition involves capturing both spatial and temporal features from video sequences to understand and classify human actions. This task can be executed in either a two-stage or one-stage configuration. In a two-stage setup, spatial features are extracted from each frame, followed by the subsequent modeling of temporal dependencies across the frames. Conversely, in a one-stage configuration, the unified architecture processes all input video frames jointly, capturing both spatial and temporal features. The subsequent descriptions provide more insight into these action recognition models.

### 04 Mar 2024 v2

@@ 3.2 Overview of vision-based seizure semiology analysis
Seizures characterized by motor manifestations are analyzed and can be automatically classified based on the specific type of visible (mainly motor) symptomatology. Some seizures are defined by unnatural elementary movements, such as those seen in myoclonic, clonic, tonic, versive, and tonic-clonic seizures [80]. The machine-based classification depends on factors such as the duration of muscle contraction, rhythmicity of movement repetition, and the muscle groups or body segments involved (e.g. asymmetric posture, flexion of the neck, abduction of both arms, and turning of eyes and head to one side). In contrast, more complex motor manifestations may sometimes seem naturalistic, and usually comprise action patterns involving movements with more complex temporal organization involving different body segments (e.g. manual and oral automatisms such as chewing, swallowing, smacking the lips and fumbling with clothes) [169, 33].
Some existing research on automated quantification of semiology is based on the hypothesis that seizures with similar motor symptoms may involve neuronal activity within the same specific brain networks, and thus could be sufficient to detect and categorize patients with specific types and brain localizations of epilepsy [132]. Vision-based approaches for semiology highlight the significance of analyzing spatio-temporal information during seizures (i.e. behavioral sequences), employing action recognition methodologies. In this domain, the discriminative power of deep learning architectures is exploited for markerless motion capture (MoCap) of body regions for isolated clinical manifestations from the body, head, face, and hands. Furthermore, deep learning is integral to seizure detection (identifying epileptic events within a continuous data stream) and seizure classification (categorizing seizures or seizure disorders).
@@
Deep learning models play a crucial role in advancing motion capture technologies, particularly for detecting regions of interest (ROI). These ROIs may encapsulate entities like patients, heads, faces, or hands. Yet these techniques go beyond mere object identification and localization, inferring facial landmarks, body keypoints, and hand landmarks from video data. While seizures, strictly speaking, may not be categorized as actions, HAR models are designed to encode robust representations of human motion that may prove relevant for seizure characterization. These methods are commonly structured to classify human actions by aggregating predictions for snippets sampled from short clips. Traditionally, a subject-level prediction is obtained by averaging all predictions made at the snippet level.
@@
Figure 1: Overview of deep learning applications for semiology analysis. A. Motion capture and skeleton-based human action recognition. B. Frame-based human action recognition. C. Feature representations of the motion exhibited by a patient during a seizure, which is used to identify if new data samples are similar to known semiologies, enabling the categorization of Epilepsy types.
@@
In the pre-deep learning era, various hand-crafted feature-based approaches, such as space-time volume-based methods, space-time interest point-based methods, and trajectory-based methods, were employed for RGB video-based detection and HAR. However, with the advent of deep learning, contemporary research in semiology has shifted to the adoption of different types of deep learning frameworks. These frameworks fall into two categories, skeleton-based and frame-based HAR. A visual representation of these frameworks is provided in Figure 1. The first, skeleton-based HAR, focuses on detecting semantic key points or movement direction/intensity initially. Subsequently, quantitative kinematic and spectral parameters are computed to detect/classify Epilepsy types with classical machine learning or deep learning methods. Here, skeleton information refers to body, facial, or hand landmarks. These approaches have been utilized by [10, 9, 110] (See Figure 1A). The second framework, a frame-based HAR, captures the dynamic variation of the body’s physical appearance from sequences of images. This approach considers both the presence of specific movements of interest (MOI) and their dynamics and biomechanical characteristics (e.g. speed/acceleration patterns, movement amplitude). Works including  [8, 120, 182, 253, 104] are example of this approach (See Figure 1B). Seizure semiology encompasses the study of the stepwise/temporal progression of a constellation of signs that are reflective of dynamic changes within connected neuronal networks. Thus, investigating a single clinical sign in isolation is generally less informative than considering all signs jointly and in the context of their usual temporal relationship. As such, multi-stream frameworks have been proposed, capable of jointly quantifying semiologic features from different body locations through fusion or hierarchical approaches (e.g. an ensemble model learning from pose and face skeleton-based streams). Hybrid networks, which fuse the benefits of both skeleton-based and frame-based approaches for seizure detection and classification, have also been proposed in [4, 103, 105]. Characteristic combinations of clinical features of different seizure types can be demonstrated to be associated with seizures arising from different brain regions at group level, although predictive power and degree of specificity/sensitivity of these varies. For some types of epilepsy, it is the motor features of the seizures that seem to most strongly predict brain localization; this has been demonstrated at the sub-lobar level in frontal lobe epilepsy [35]. Such seizure types with prominent and potentially discriminating motor patterns lend themselves well to analysis using computer vision-based approaches. A visual example of these motions is illustrated in Figure 1C. The utilization of these feature representations contributes to the development of robust models capable of discerning and categorizing diverse seizure types. Further details on recent case studies of vision-based approaches in Epilepsy are discussed in Section 4. The following subsections provide details of the specific techniques used by each human action recognition framework.
@@ 3.3 Skeleton-based human action recognition
In skeleton-based action recognition, the primary focus is on explicitly tracking and analyzing the movement and positions of skeletal joints from the body, face, or hands, that inherently contain spatio-temporal information. Skeleton data offers numerous advantages for HAR including a simple yet informative representation, scale invariance, and robustness to variations in clothing textures and backgrounds. These approaches are often fast since the human pose representation is very compact. However, they lack appearance and detailed shape information and the underlying skeleton representation can be considered a noisy representation [211]. Another drawback is the propagation of errors from the motion capture to the action recognition stage, a factor highly dependent on the quality of the spatio-temporal feature extraction of the initial stage. Joint information is vulnerable to conditions where joint estimation is performed erroneously, which is not uncommon due to the frequent occlusions that occur in seizure videos. Figure 2 illustrates these stages.
@@ 3.3.1 Markerless motion capture: keypoint detection
Human pose estimation: Human pose estimation involves localizing anatomical keypoints, and is widely studied in computer vision and has broad applications (e.g. motion analysis, healthcare, virtual reality). Although camera-based methods for human pose estimation have been studied extensively, in-bed pose analysis comes with specific challenges including joints being occluded, blur, and low-light conditions. Traditional pose estimation models are broadly categorized into two groups: in the first approach, a human bounding box is initially detected, and then keypoints indicating joint locations for each person are estimated (known as top-down methods, e.g. AlphaPose [69], ViTPose [247]). This can lead to errors as the predicted bounding box in the first step may not include the required joints for the second step. In the second approach, all keypoints are detected in the first stage and they are assigned to each person in an image in a second stage (bottom-up methods, or box-free human detection, e.g. OpenPose [37]). Human pose estimation has also been investigated in 3D space to predict the locations of body joints. For instance, a 3D human pose estimation algorithm was developed by taking into account spatial and temporal information [215]. In addition to estimating the 3D pose, some methods also recover a 3D human mesh from images or videos [150, 176]. This field has attracted much interest in recent years since it is used to provide extensive 3D structure information related to the human body.
@@
Hand landmark estimation: Hand pose estimation refers to methods used to detect hand landmarks or keypoints. This is of interest given the rich and subtle repertoire of abnormal hand postures and movements observed in epileptic seizures, in conjunction with the rapid technical progress in automated methods for hand detection in video [78, 209]. Methods that can localize hand joints share several properties with human body pose estimation, and many approaches proposed for the human body can be adapted for hand pose estimation. Similar to body pose, hand pose can also be estimated in 2D [201, 234] or 3D [173, 50] space. Image generation and synthetic data have also become popular in hand pose estimation. Examples of recently introduced synthetic data are  [126], which includes multi-view hand data, and [142], which features 1 million photo-realistic images of hands interacting with various objects, backgrounds, and textures. For more information about 2D/3D or hybrid hand pose models, refer to [49, 170].
@@ 3.3.2 Action recognition models (skeleton-based HAR)
CNN-based methods: This approach utilizes the advantages of CNN-based object recognition in a 2D space (image space). It maps a 3D representation of the 3D human skeletons into a 2D array (possibly capturing the spatial relationships between the skeleton joints) to learn spatio-temporal skeleton features. Examples of recent works using CNN-based spatio-temporal approaches are CNN fusion [144], and PoseConv3D [65]. A CNN fusion model for skeletal action recognition is proposed in [144], which uses spatio-temporal information including skeletal trajectory shape images as well as skeletal pose image sequences. PoseConv3D [65] used the 3D-CNN to capture the spatio-temporal dynamics of skeleton sequences; where the input to the 3D-CNN backbone are 3D heatmap volumes. The pseudo heatmaps for joints and limbs that are generated provide rich information as inputs for 3D-CNNs. More information about convolutional neural network-based action can be found in [254].
@@
GNN-based methods: Early approaches to skeleton-based HAR employed RNNs and CNNs, but recent advancements have surpassed them with the introduction of Graph Convolutional Networks (GCNs). Leveraging the expressive power of graph structures, GCNs efficiently represent non-Euclidean data, allowing them to capture both spatial (intra-frame) and temporal (inter-frame) information. These learning models have been adopted for various tasks including skeleton-based human action recognition [199, 172] and facial landmark-based emotion recognition [168]. The spatio-temporal graph (ST-GCN) [250], the first work utilizing GCNs for HAR, is constructed such that the graph nodes represent the joint and the edges encode natural connections in both human body and time. Multiple layers of spatio-temporal graph convolution operations are applied to the input data, generating higher-level feature maps on the graph, subsequently classifying the input into the corresponding action category. Followed by the success of ST-GCN, actional-structural graphs have been proposed to capture both actional links and structural links over time [143]. Another work involves the use of a graph network to perform spatial reasoning and temporal stack learning (SR-TSL) for skeleton-based HAR, comprising a spatial reasoning network (SRN) and a temporal stack learning network (TSLN) [200]. More recent approaches include ST-GCN++ [64] and Efficient GCN [207].
Transformer-based methods: Vision transformers are emerging as a powerful tool to solve computer vision problems. Recent techniques have also proven the efficacy of transformers beyond the image domain, extending to diverse video-related tasks, including action recognition [17].
@@
The architecture of transformers, featuring a self-attention mechanism, proves useful for estimating temporal relationships and motion dynamics encoded in joint positions. Some examples of these approaches include PoseConv3D [65], SkeletonMAE [244] and InfoGCN++ [51]. PoseConv3D [65] addresses limitations in current Skeleton-based Human Activity Recognition (HAR) methods by combining 3D CNN and a transformer-based approach for feature extraction, modeling temporal dependencies, and handling the sparsity of skeleton data. SkeletonMAE [244] focuses on capturing structural information and temporal dependencies in skeleton data by representing it as a graph. It introduces masked autoencoders to learn meaningful representations by reconstructing the original skeleton sequences with masked nodes. Furthermore, InfoGCN++ [51], leverages graph convolutional networks (GCNs) as part of the transformer network to learn informative representations by predicting future frames in the skeleton sequence. These transformer-based methods incorporate temporal information through a joint learning strategy, resulting in improved performance in action recognition tasks.
3.4 Frame-based HAR modality
@@
In frame-based action recognition, the focus is on analyzing regions or patches within sequences of images. This approach does not rely on explicit skeletal data but rather on the visual information captured in the frames, including the spatial layout and temporal changes in these regions [211]. This framework is illustrated in Figure 3. As these approaches directly operate on RGB videos, there is a possibility of privacy leakage of sensitive patient data from videos. Moreover, obtaining consent from patients to share their raw RGB video data for inter-cohort validation studies and developing approaches that generalize to large populations becomes challenging.
@@
Figure 3: Overview of frame-based human action recognition. A. Motion capture: Region of interest detection such as the body, face, or hands. B. Spatio-temporal representations are learned from sequential frames. CNN-RNN-based, 3D-CNN-based and Transformer-based methods are all commonly used with this type of approach.
@@ 3.4.1 Markerless motion capture: region of interest detection
3.4.2 Action recognition models (frame-based HAR)
Frame-based human action recognition involves capturing both spatial and temporal features from video sequences to understand and classify human actions. This task can be executed in either a two-stage or one-stage configuration. In a two-stage setup, spatial features are extracted from each frame, followed by the subsequent modeling of temporal dependencies across the frames. Conversely, in a one-stage configuration, the unified architecture processes all input video frames jointly, capturing both spatial and temporal features. The subsequent descriptions provide more insight into these action recognition models.